# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v2

@@ 5 ICE-INTERN: Bilingual Financial Large Language Model
| Specific Task | Metric | Lang | Data | Test | Cover | ||||||
| [1] | [2] | [3] | [4] | [5] | [6] | ||||||
| Sentiment Analysis | FinSA | F1 Accuracy | zh | FE | 2,020 | ||||||
| StockB | 1,962 | ||||||||||
| CFPB | 970 | ||||||||||
| CFiQA-SA | 233 | ||||||||||
| en | FPB | 970 | |||||||||
| FiQA-SA | 235 | ||||||||||
@@
F1 Accuracy | zh | Corpus AFQMC | 10,000 4,316 | ||||||
@@
F1 Accuracy | zh | FinevalF | 222 | ||||||
| Stock Prediction | FinSP | F1 Accuracy MCC | zh | StcokA | 1,477 | ||||||
| CACL18 | 511 | ||||||||||
| CBigData18 | 159 | ||||||||||
| CCIKM18 | 86 | ||||||||||
| en | ACL18 | 3,720 | |||||||||
| BigData18 | 1,472 | ||||||||||
| CIKM18 | 1,143 | ||||||||||
@@
F1 Accuracy MCC | en | German Australian | 200 139 | ||||||
@@
F1 Accuracy | en | FOMC | 496 | ||||||
| Question Answering | FinQA | EM Accuracy | zh | QA | 2,469 | ||||||
| CEnQA | 133 | ||||||||||
| CConFinQA | 237 | ||||||||||
| en | EnQA | 1,147 | |||||||||
| ConFinQA | 1,490 | ||||||||||
@@
F1 Entity F1 | zh en en | CNER NER FINER-ORD | 337 98 1,075 | ||||||
| Event Detection | FinED | F1 Recall Precision | zh | 19CCKS | 2,936 | ||||||
| 20CCKS | 9,159 | ||||||||||
| 21CCKS | 1,400 | ||||||||||
| 22CCKS | 11,829 | ||||||||||
@@
Rouge BERTScore BARTScore | zh en en | NA ECTSUM EDTSUM | 3,600 495 2000 | ||||||
| Open Source | |||||||||||
| Task Total | 14 | 5 | 6 | 8 | 4 | 1 | 1 | ||||
| Data Total | 95k | 43k | 4k | 11k | 4k | 5k | 0.2k | ||||
Upon evaluating performance of various backbone LLMs on ICE-FLARE, we explore InternLM-7B to bulid ICE-INTERN-7B with the ICE-FIND. To assess the impact of different data types on fine-tuning ICE-INTERN, we gradually incorporate the DLC, DLE, DTT and DTE datasets during the fine-tuning process to create variants of ICE-INTERN named ICE-INTERN-dlc-7B, ICE-INTERN-dle-7B, ICE-INTERN-dtt-7B, and ICE-INTERN-full-7B, respectively. We use the same set of fine-tuning hyper-parameters for all models, which are described below. To improve the model’s capability in handling long text, we employ an advanced fine-tuning technique called QLoRA Hu et al. (2021) to enhance the training process, leading to optimized memory usage and faster training times. Specifically, the training sequences are uniformly cut to a length of 2048 tokens. Then, we use AdamW optimizer with an initial learning rate 5e-5 on weight decay 1e-5 for all parameters. Additionally, the warmup steps are set to 1% of the total training steps. Finally, we fine-tune model total 1 epoch in batch size of 24 on 8 A100 40GB GPUs. Due to the growing volume of fine-tuning data and constraints on our computational resources, we have opted not to proceed with additional fine-tuning of ICE-INTERN using larger parameters.
6 ICE-FLARE: Bilingual Financial Evaluation Benchmark
| Model | ZH-CLS | ZH-PRE | ZH-EXT | ZH-GEN | ZH-TRA | ZH.Avg | EN-CLS | EN-PRE | EN-EXT | EN-REA | EN-OFT | EN.Avg | Bi.Avg | CLS.Avg | PRE.Avg | EXT.Avg |
@@ 6 ICE-FLARE: Bilingual Financial Evaluation Benchmark
| CFGPT-sft-7B | 0.211 | 0.050 | 0.011 | -2.017 | 0.272 | -0.295 | 0.242 | 0.224 | 0.000 | 0.000 | -1.265 | -0.160 | -0.227 | 0.227 | 0.137 | 0.006 |
| DISC-FinLLM-13B | 0.290 | 0.315 | 0.022 | -1.858 | 0.290 | -0.188 | 0.446 | 0.236 | 0.039 | 0.002 | -1.158 | -0.087 | -0.138 | 0.368 | 0.276 | 0.031 |
| FinMA | 0.229 | 0.000 | 0.016 | -2.009 | 0.250 | -0.303 | 0.371 | 0.233 | 0.392 | 0.000 | -1.187 | -0.038 | -0.171 | 0.300 | 0.117 | 0.204 |
| ICE-INTERN-dlc-7B | 0.598 | 0.340 | 0.020 | -1.762 | 0.337 | -0.093 | 0.331 | 0.282 | 0.000 | 0.000 | -1.227 | -0.123 | -0.108 | 0.465 | 0.311 | 0.010 |
| ICE-INTERN-dle-7B | 0.607 | 0.637 | 0.569 | -1.024 | 0.327 | 0.223 | 0.334 | 0.275 | 0.006 | 0.000 | -1.203 | -0.118 | 0.053 | 0.471 | 0.456 | 0.288 |
| ICE-INTERN-dtt-7B | 0.607 | 0.630 | 0.566 | -1.024 | 0.448 | 0.245 | 0.357 | 0.268 | 0.000 | 0.000 | -1.182 | -0.111 | 0.067 | 0.482 | 0.449 | 0.283 |
| ICE-INTERN-full-7B | 0.631 | 0.617 | 0.567 | -0.972 | 0.409 | 0.250 | 0.494 | 0.250 | 0.362 | 0.000 | -1.193 | -0.017 | 0.117 | 0.563 | 0.434 | 0.465 |
Based on ICE-FIND, we design our bilingual multi-task evaluation benchmark, ICE-FLARE. The existing benchmarks ([1] FLARE Xie et al. (2023b), [2] CFBenchmark Lei et al. (2023), [3] FinCUGE Chen et al. (2023), [4] FLUE Shah et al. (2022), [5] Fineval Zhang et al. (2023b), [6] CGCE Zhang et al. (2023c)) compared for ICE-FLARE are listed in Table 2. Unfortunately, they are all designed exclusively for either Chinese or English. ICE-FLARE has highlighted its superiority over FLUE in English through FLARE benchmark Xie et al. (2023b). From Table 2, it is evident that ICE-FLARE introduce bilingual capabilities, significantly surpasses these benchmarks on task diversity and data scalability. Despite FinCUGE having a higher number of financial evaluation tasks, the difference in dataset size compared to ICE-FLARE is nearly 8 times. These public benchmarks have provided valuable references and inspiration for constructing the test set, making ICE-FLARE more comprehensive in terms of task coverage, metric diversity, and sample quantity. However, they still face challenges in evaluating the bilingual capabilities in Chinese-English and severely lack DOF tasks (e.g., FinDC) to test the robustness of the models. In conclusion, ICE-FLARE, which offers a bilingual evaluation benchmark and a more comprehensive assessment of diversity tasks, is necessary for improving performance and linguistic capabilities of FinLLMs.
7 Experiments
7.1 Evaluation Setup
In addition to the FinLLMs we mentioned earlier, such as CFGPT, DISC-FinLLM, FinMA, we also select the representative and outstanding LLMs for evaluating performance. 1) ChatGPT (OpenAI Brown et al. (2020)): ChatGPT-3.5-Trubo and GPT4, developed by OpenAI, are widely recognized LLMs with versatile functionalities. 2) LLaMa Touvron et al. (2023): LLaMa and advanced LlaMa2 are powerful LLMs developed by Meta AI, with parameter sizes ranging from 7 billion to 650 billion. 3) Baichuan Yang et al. (2023a): It is a LLM developed by Baichuan-inc for Chinese and English NLP tasks, available in Baichuan-7B and Baichuan-7B-Chat variants. 4) ChatGLM Du et al. (2021): It is an advanced bilingual LLM developed by Tsinghua University and ZhiPu AI, with versions like ChatGLM2-6B and ChatGLM3-6B. 5) BLOOM Muennighoff et al. (2022): It developed by the BigScience team, is a versatile LLM with parameter sizes ranging from 560M to 176B, including Bloomz-7B1. 6) InternLM Team (2023): InternLM is a language model developed by SenseTime, featuring different parameter configurations such as InternLM-7B. 7) Qwen Bai et al. (2023): It is a series of advanced LLMs introduced by Alibaba Cloud, including Qwen-7B for conversational capabilities in Chinese.
7.2 Results
(1) Benchmark Evaluation Performance The bilingual ICE-FLARE benchmark results are shown in Table 3. As shown in Table 3, the performance of ICE-INTERN and its variants in monolingual settings significantly favors Chinese over English, both across various tasks and in overall average performance. In English, the overall average performance of fully fine-tuned ICE-INTERN-full is slightly weaker than that of ChatGPT. However, across individual tasks, except for ChatGPT, ICE-INTERN and its variants perform best in the CLS and PRE NLP tasks, while their performance in the REA and OFT NLP tasks is not particularly remarkable, sometimes even weaker than other models such as FinMA and Bloomz-7B1. This indicates that full fine-tuning ICE-INTERN-full leads to a substantial overall improvement, offsetting some subpar task performances. Furthermore, we observed that apart from GPT4, fully fine-tuned ICE-INTERN-full demonstrates superior performance in bilingual tasks, particularly evident in the PRE and EXT tasks. (2) Model Balanced Performance As shown in Figure 4, A demonstrates that ICE-INTERN excels in specific tasks, particularly those involving DLC, DLE, and DTT data types, showcasing strong Chinese language capabilities. However, in D-E, it slightly lags behind GPT4 in English and DOF tasks, a distinction notably evident in F. Within G, ICE-INTERN excels in 10 Chinese-English bilingual NLP tasks, except in English reasoning and DOF tasks. H further confirms ICE-INTERN ’s overall superiority in handling tasks in both English and Chinese, particularly excelling in Chinese.
We also draw the conclusions from the detailed results in Appendix Table 11 .
(3) Model Detail Performance Among the total of 63 specific-task metrics, ICE-INTERN variants achieves optimal performance in approximately 68%. Among them, ICE-INTERN-Full, which fine-tuned on the entire dataset, exhibits the highest percentage of all tasks with the best performance. This fully demonstrates the exceptional superiority of ICE-INTERN and highlights the importance of domain instruction fine-tuning in improving the performance of FinLLMs. InternLM overall outperforms LLaMA2, which fully validates our choice of using InternLM instead of LLaMA2 as the backbone model used in similar work Yang et al. (2023c); Wu et al. (2023); Xie et al. (2023b). We observe that that DLC has little impact for ICE-INTERN-dlc-7B. New introduction of DLE further improves the performance of ICE-INTERN-dle-7B significantly. Conversely, introducing DTT for fine-tuning ICE-INTERN-dtt-7B does not result in an improvement and instead leads to decrease. Finally, incorporating DTE for fine-tuning ICE-INTERN-Full-7B demonstrates a substantial improvement. Our discoveries align with prior research Chen et al. (2023); Xie et al. (2023b), illustrating the efficacy of the InternLM backbone in DLC. Additionally, diverse DLE data types encompassing comprehension and reasoning tasks prove essential for optimizing performance. Moreover, our results underscore the detrimental impact of poor-quality translation datasets on fine-tuned model efficacy. These insights pave the way for innovative advancements in LLM research within the finance domain. (4) Bilingual Proficienc Analysis In the cross-lingual tasks of FinHC, FinSA, and FinSP, the ICE-INTERN-Full model exhibits performance improvements ranging from approximately 1% to 10% when incorporating the corresponding translation data in fine-tuning LLMs for the same task. These findings suggest that the impact of supplementing the corresponding translation data on fine-tuning results is not as pronounced for shorter-text tasks (such as FinSA and FinHC) in DLC. However, when dealing with the long-text (including time-series data) task of FinSP in DLC, incorporating the corresponding translation data can significantly enhance the results. This is because short-text specific tasks involves relatively simple grammar and structure of the financial descriptions, and our ICE-INTERN models can already utilize sufficient information in Chinese for answer reasoning. However, when dealing with longer and more complex texts such as those containing time, historical stock prices, and tweets, our ICE-INTERN models benefit from the inclusion of English data, which enhances its ability to reason with contextual information. These observations highlight the importance of considering the length and complexity of the text when determining the effectiveness of incorporating translation data for fine-tuning LLMs in cross-lingual tasks. (5) Defect Case Analysis Despite the strengths exhibited by our ICE-INTERN models on known tasks, they under perform relative to GPT4 on unseen tasks like FinDC, FINER, and FinST in DOF data types, which have not been fine-tuned. This performance gap suggests that a more diverse set of domain-specific tasks is needed for effective fine-tuning. Specifically, in FinST tasks, we utilize a sophisticated prompt design that involves generating label sequences directly. However, our fine-tuned models consistently struggle to produce outputs in the desired formats. Even models like larger-parameter ChatGPT-3.5 and GPT4, which demonstrate some capability for token labeling in FinER, face difficulties in generating labels and summaries for ECTSUM and EDTSUM in FinST, especially when confronted with longer contextual information. Our conclusion aligns with PIXIU, suggesting that for unfamiliar tasks within current FinLLMs, it will be imperative to furnish more extensive domain-specific tasks and instructional data in the future to further enhance model performance.
7.3 Practical Examples

### v3

@@ 3.3. ICE-FLARE: Bilingual Financial Evaluation Benchmark
| Specific Task | Metric Language | Data | Test | Chinese | English | ||||||||||
| [a] | [b] | [c] | [d] | [e] | [f] | [g] | [h] | [i] | [j] | ||||||
| Sentiment Analysis | FinSA | F1 Accuracy | zh | FE | 2,020 | ||||||||||
| StockB | 1,962 | ||||||||||||||
| CFPB | 970 | ||||||||||||||
| CFiQA-SA | 233 | ||||||||||||||
| en | FPB | 970 | |||||||||||||
| FiQA-SA | 235 | ||||||||||||||
@@
F1 Accuracy | zh | Corpus AFQMC | 10,000 4,316 | ||||||||||
@@
F1 Accuracy | zh | FinevalF | 222 | ||||||||||
| Stock Prediction | FinSP | F1 Accuracy MCC | zh | StcokA | 1,477 | ||||||||||
| CACL18 | 511 | ||||||||||||||
| CBigData18 | 159 | ||||||||||||||
| CCIKM18 | 86 | ||||||||||||||
| en | ACL18 | 3,720 | |||||||||||||
| BigData18 | 1,472 | ||||||||||||||
| CIKM18 | 1,143 | ||||||||||||||
@@
F1 Accuracy MCC | en | German Australian | 200 139 | ||||||||||
@@
F1 Accuracy | en | FOMC | 496 | ||||||||||
| Question Answering | FinQA | EM Accuracy | zh | QA | 2,469 | ||||||||||
| CEnQA | 133 | ||||||||||||||
| CConFinQA | 237 | ||||||||||||||
| en | EnQA | 1,147 | |||||||||||||
| ConFinQA | 1,490 | ||||||||||||||
@@
F1 Entity F1 | zh en en | CNER NER FINER-ORD | 337 98 1,075 | ||||||||||
| Event Detection | FinED | F1 Recall Precision | zh | 19CCKS | 2,936 | ||||||||||
| 20CCKS | 9,159 | ||||||||||||||
| 21CCKS | 1,400 | ||||||||||||||
| 22CCKS | 11,829 | ||||||||||||||
@@
Rouge BERTScore BARTScore | zh en en | NA ECTSUM EDTSUM | 3,600 495 2000 | ||||||||||
| New Task Total | 0 | 2 | 0 | 0 | 0 | 1 | 0 | 0 | 1 | 1 | |||||
| Open Source | |||||||||||||||
| Task Total | 14 | 6 | 8 | 1 | 1 | 1 | 4 | 8 | 1 | 2 | 8 | ||||
| Data Total | 95k | 4k | 11k | 5k | 0.2k | 9.9k | 4k | 43k | 10k | 5.5k | 54k | ||||
4. Experiments
In addition to the FinLLMs such as CFGPT-sft (CFGPT), DISC-FinLLM (DISCFIN), FinMA, we also select the most representative baseline LLMs for performance evaluation on ICE-FLARE.
GPT-4 (Achiam et al., 2023): An OpenAI-developed powerful LLM with around 1T parameters, capable of handling diverse modalities including images and videos.
ChatGPT (Brown et al., 2020): It also known as ChatGPT-3.5-Turbo-1106 with 175B parameters from OpenAI, is a widely-known LLM in the NLP domain.
LLaMa (Touvron et al., 2023): LLaMa and advanced LlaMa2 are powerful LLMs developed by Meta AI, with parameter sizes ranging from 7 billion to 650 billion.
Baichuan (Yang et al., 2023c): This LLM developed by Baichuan-inc for Chinese and English NLP tasks, available in Baichuan-7B and Baichuan-7B-Chat variants.
ChatGLM (Du et al., 2021): It is an advanced bilingual LLM developed by Tsinghua University and ZhiPu AI, with variants like ChatGLM2-6B and ChatGLM3-6B.
BLOOM (Muennighoff et al., 2022): Developed by the BigScience team, it is a versatile LLM with parameter sizes ranging from 560M to 176B, including Bloomz-7B1.
InternLM (Team, 2023): InternLM is a language model developed by SenseTime, featuring different parameter configurations such as InternLM-7B.
Qwen (Bai et al., 2023): It is a set of advanced LLMs introduced by Alibaba Cloud, including Qwen-7B for conversational capabilities in Chinese.
All results from all LLMs are conducted in the zero-shot setting (Li et al., 2023b) with automatic evaluation method.
| Model | ZH-CLS | ZH-PRE | ZH-EXT | ZH-GEN | ZH-TRA | ZH.Avg | EN-CLS | EN-PRE | EN-EXT | EN-REA | EN-OFT | EN.Avg | Bilingual.Avg | CLS.Avg | PRE.Avg | EXT.Avg |
@@ 4. Experiments
| CFGPT-7B | 0.211 | 0.050 | 0.011 | -2.017 | 0.272 | -0.295 | 0.242 | 0.224 | 0.000 | 0.000 | -1.265 | -0.160 | -0.227 | 0.227 | 0.137 | 0.006 |
| DISCFin-13B | 0.290 | 0.315 | 0.022 | -1.858 | 0.290 | -0.188 | 0.446 | 0.236 | 0.039 | 0.002 | -1.158 | -0.087 | -0.138 | 0.368 | 0.276 | 0.031 |
| FinMA-7B | 0.229 | 0.000 | 0.016 | -2.009 | 0.250 | -0.303 | 0.371 | 0.233 | 0.392 | 0.000 | -1.187 | -0.038 | -0.171 | 0.300 | 0.117 | 0.204 |
| ICEdlc-7B | 0.598 | 0.340 | 0.020 | -1.762 | 0.337 | -0.093 | 0.331 | 0.282 | 0.000 | 0.000 | -1.227 | -0.123 | -0.108 | 0.465 | 0.311 | 0.010 |
| ICEdle-7B | 0.607 | 0.637 | 0.569 | -1.024 | 0.327 | 0.223 | 0.334 | 0.275 | 0.006 | 0.000 | -1.203 | -0.118 | 0.053 | 0.471 | 0.456 | 0.288 |
| ICEdtt-7B | 0.607 | 0.630 | 0.566 | -1.024 | 0.448 | 0.245 | 0.357 | 0.268 | 0.000 | 0.000 | -1.182 | -0.111 | 0.067 | 0.482 | 0.449 | 0.283 |
| ICEfull-7B | 0.631 | 0.617 | 0.567 | -0.972 | 0.409 | 0.250 | 0.494 | 0.250 | 0.362 | 0.000 | -1.193 | -0.017 | 0.117 | 0.563 | 0.434 | 0.465 |
4.1. Results
The results for ICE-FLARE benchmark are shown in Table 4 and Figure 2. More detailed results can be found in Appendix Table 11.
4.1.1. Overall Performance
As shown in Table 4, ICE-INTERN variants outperform all baseline LLMs in Chinese tasks, even surpassing larger LLMs like GPT-4. ICE-INTERN falls slightly behind GPT-4 in average performance across bilingual tasks but surpasses all other LLMs. It excels in CLS and PRE bilingual tasks, outperforming all models. However, its performance in REA and OFT tasks is not particularly remarkable, sometimes even weaker than other LLMs such as FinMA-7B and Bloomz-7B1. This clearly demonstrates the absolute superiority of ICEfull-7B in the Chinese-English financial bilingual domain, emphasizing the importance of instruction fine-tuning over the InternLM backbone model in optimizing the financial performance of various LLMs. ICE-INTERN variants like ICEdle-7B and ICEdtt-7B, which are not fine-tuned with English datasets, thus performing slightly below GPT-4 due to its larger parameters and emphasis on English. This suggests that there is further optimization potential in future model iterations by exploring the introduction of more effective English data. Referring to Table 11, ICE-INTERN variants achieve the best performance in 68.25% (excluding ChatGPT and GPT-4) and 58.73% of the 63 specific-task metrics. In terms of overall best results, the fully fine-tuned ICEfull-7B outperforms all baseline models, notably outperforming both ChatGPT and even surpassing larger models like GPT-4.
4.1.2. Language Disparity
Table 4 highlights significant language discrepancies in how current LLMs, such as well-known ChatGPT, tackle Chinese-English bilingual financial NLP tasks. The findings suggest that these LLMs, including the Chinese-focused Qwen-7B, excel predominantly in English but exhibit slight weaknesses in Chinese. Our proposed ICE-INTERN and its variations distinctly prioritize Chinese proficiency over English, both in individual tasks and across overall average metrics. This indicates that addressing language disparities in LLMs requires dedicated efforts to construct and optimize high-quality instructions with multilingual characteristics. Further referencing Figure 2, A-C indicates that ICEfull-7B excels in specific tasks involving DLC, DLE, and DTT data types, showcasing strong Chinese language capabilities. However, in D-E, it slightly lags behind GPT-4 in the FinER, FinQA, FinTS, FinCC tasks, with the performance gap being particularly pronounced in F. In G, ICEfull-7B excels in 10 Chinese-English bilingual NLP tasks, except for English extraction (EN-EXT), reasoning (EN-REA) and data-out-of-training (EN-DOF) tasks. H further confirms ICEfull-7B’s overall superiority in handling tasks in both English and Chinese, with particular excellence in Chinese.
4.1.3. Generalization Ability
Despite the strengths exhibited by ICE-INTERN and its variants on known specific tasks, they lags behind GPT-4 in unknown tasks like FinDC, FINER, and FinST within the unfine-tuned DOF data type. This performance gap indicates the need for a wider variety of domain-specific financial tasks for effective model fine-tuning. Specifically, we employ complex prompt designs involving directly generating label sequences in the FinST task. However, our fine-tuned ICE-INTERN models consistently encounter difficulties in producing outputs in the desired formats. Even LLMs such as larger-parameter ChatGPT and GPT-4, which demonstrate some capability for token labeling in FinER, encounter challenges in generating labels and summaries for both the ECTSUM and EDTSUM datasets in FinST, especially when dealing with longer financial contextual information. Our findings are in line with the conclusions put forth in existing studies (Xie et al., 2023), suggesting that for unfamiliar financial tasks within current FinLLMs, it will be imperative to provide more extensive domain-specific tasks and instruction data in the future to further enhance model performance.
4.1.4. Ablation study
This ablation study primarily analyzes the performance impact of variant models on different datasets (DLC, DLE, DTT, DTE, and DOT). As shown in Table 11, DLC has minimal impact on the performance of ICEdlc-7B, while fine-tuning with DLE significantly improves the performance of ICEdle-7B. Conversely, fine-tuning ICEdle-7B with DTT does not enhance the performance of ICEdtt-7B and may even lead to a decrease. Finally, full fine-tuning of ICEFull-7B with DTE shows a significant performance improvement. This indicates that InternLM, as the backbone model, is already capable of handling most simple financial classification tasks present in DLC data types, making fine-tuning on such data less consequential. Clearly, it is essential to incorporate a variety of DLE data types in the instruction data to enhance model performance for comprehension and inference in more intricate tasks. Furthermore, we find that in the cross-lingual financial classification tasks in FinHC and FinSA, the ICEdtt-7B model shows performance improvements of approximately 1% to 10% compared to ICEdle-7B when incorporating Chinese translation datasets. However, for the complex FinSP task, the performance exhibits a mix of improvements and declines. This is because tasks like FinHC and FinSA involve relatively simple financial description syntax and structure, allowing ICEdtt-7B to integrate translated data to learn sufficient information for answer reasoning. However, for tasks like FinSP, which entail longer and more complex financial text including historical stock prices and tweets, data augmentation methods did not yield the anticipated model performance enhancements. This emphasizes the need to consider text length and task complexity when deciding on incorporating translation data during LLM fine-tuning.
4.2. Practical Examples